For a gaseous reaction, standard conditions are 29
Answer #1

Q = PNH3^2 / PN2 x PH2 ^3

Q = 0.750 ^2 / 0.500 x (0.350)^3

Q = 26.24

\DeltaGo =-72.6 kJ / mol

T = 298 K

R = 8.314 x 10^-3 kJ / mol K

\DeltaG = \DeltaGo   + R T ln Q

         = -72.6 + 8.314 x 10^-3 x 298 x ln 26.24

        = -64.5 kJ / mol

\DeltaG = - 64.5 kJ / mol